# The commutativity of cofinal types

### v1

@@ 3. Proof of the main result and some corollaries
For any non-principal ultrafilter $U,V$ over $\omega$, $V\cdot U,\ U\cdot V\geq_{T}(U^{*}\cap V^{*})^{\omega}$
@@
We have that $U^{*}\cap V^{*}\subseteq V^{*}$ and $\mathrm{fin}\subseteq V^{*}$. We conclude that $(U^{*}\cap V^{*})\otimes\mathrm{fin}\subseteq(V\cdot V)^{*}$. By Corollary 2.31, $(U^{*}\cap V^{*})\otimes\mathrm{fin}$ is deterministic, hence
$U^{*}\cap V^{*}\leq_{T}(U^{*}\cap V^{*})\otimes\mathrm{fin}\leq_{T}V\cdot V$
By Milovich’s Corollary 1.8 and Theorem 1.7
$U\cdot V\equiv_{T}U\cdot(V\cdot V)\equiv_{T}U\times(V\cdot V)^{\omega}\geq_{T}(U^{*}\cap V^{*})^{\omega}.$
The proof that $V\cdot U\geq_{T}(U^{*}\cap V^{*})^{\omega}$ is symmetric. ∎
@@
Suppose that $U,V$ are ultrafilters on $\omega$, then
$U\cdot V\equiv_{T}U\times V\times(U\cap V)^{\omega}.$
Hence $U\cdot V\equiv_{T}V\cdot U$.
@@
By the previous corollary, $(U\cap V)^{\omega}\equiv_{T}(U^{*}\cap V^{*})^{\omega}\leq_{T}U\cdot V$. Since $U\times V\times(U\cap V)^{\omega}$ is the least upper bound in the Tukey order of $U,V,(U\cap V)^{\omega}$, we have
$U\times V\times(U\cap V)^{\omega}\leq_{T}U\cdot V$
For the other direction, recall that by Theorem 2.13 $V$ has the $(U\cap V)^{*}$-p.i.p and therefore by Proposition 2.5
$U\cdot V\equiv_{T}U\times V^{\omega}\leq_{T}U\times V\times(U^{*}\cap V^{*})^{\omega}\equiv_{T}U\times V\times(U\cap V)^{\omega}.$
Recall that for an ultrafilter $U$, $Ch(U)$ denoted the minimal size of a cofinal subset of $U$. It is clear that if $U\equiv_{T}U^{\prime}$ then $Ch(U)=Ch(U^{\prime})$. As a corollary of theorem 3.2, we get:
Corollary 3.3.
For any two ultrafilters, $Ch(U\cdot V)=Ch(V\cdot U)$.
It is possible to prove by induction now that the product of $n$-ultrafilters commute, but we would like to get the exact cofinal type of such product. We will need the following facts:
Fact 3.4.
  1. (1)
    $U_{1}\cdot U_{2}\cdot...\cdot U_{n}\equiv_{T}U_{1}\times(U_{2})^{\omega}...\times(U_{n})^{\omega}$.
  2. (2)
    $(\mathbb{P}^{\omega})^{\omega}\simeq\mathbb{P}^{\omega}$.
@@
$(1)$ is just by induction using Milovich Theorem 1.7. For $(2)$, just decompose $\omega$ into infinitely many infinite sets. See for example [3, Fact 2.4]. ∎
Corollary 3.5.
For every ultrafilters $U_{1},...,U_{n}$,
$U_{1}\cdot U_{2}\cdot...\cdot U_{n}\equiv_{T}U_{1}\times U_{2}\times...\times U_{n}\times(U_{1}\cap...\cap U_{n})^{\omega}$
@@
Following the proof of Theorem 3.2, $U_{1}\cap...\cap U_{n}\subseteq U_{n}$ and we can prove similarly that $(U_{1}\cap....\cap U_{n})^{\omega}\leq_{T}U_{n}\cdot U_{n}\leq_{T}U_{1}\cdot...\cdot U_{n}$. We conclude that
$U_{1}\times...\times U_{n}\times(U_{1}\cap...\cap U_{n})^{\omega}\leq_{T}U_{1}\cdot...\cdot U_{n}$
In the other direction, each $U_{i}$ has the $(U_{1}\cap...\cap U_{n})^{*}$-p.i.p and therefore
$U_{1}\cdot...\cdot U_{n}\equiv_{T}U_{1}\times(U_{2})^{\omega}\times(U_{3})^{\omega}\times...\times(U_{n})^{\omega}\leq_{T}$
$\leq_{T}U_{1}\times U_{2}\times(U_{1}\cap....\cap U_{n})^{\omega}\times U_{3}\times(U_{1}\cap...\cap U_{n})^{\omega}\times...\times U_{n}\times(U_{1}\cap...\cap U_{n})^{\omega}$
$\leq_{T}U_{1}\times...\times U_{n}\times((U_{1}\cap...\cap U_{n})^{\omega})^{\omega}\equiv_{T}U_{1}\times....\times U_{n}\times(U_{1}\cap...\cap U_{n})^{\omega}$
Corollary 3.6.
Suppose that $U\cdot U\equiv_{T}U$ then for every $V\geq_{T}U$, $V\cdot V\equiv_{T}V$. Namely the class of ultrafilters $U$ such that $U\cdot U\equiv_{T}U$ is upward closed and the class of ultrafilters $V$ such that $V\cdot V>_{T}V$ is downward closed.
Proof.
Otherwise, $V<_{T}V\cdot V$ and
$U\cdot V\equiv_{T}U\times V^{\omega}\equiv_{T}V^{\omega}\equiv_{T}V\cdot V>_{T}V\equiv_{T}V\times U\equiv_{T}V\times U^{\omega}\equiv_{T}V\cdot U,$
contradiction. ∎
Corollary 3.7.
For every $\alpha$, $\mathrm{fin}^{\otimes\alpha}$ is deterministic.
Proof.
The base case $\mathrm{fin}^{\otimes 1}=\mathrm{fin}$ which is deterministic. Suppose this is true for $\alpha$, then $\mathrm{fin}^{\otimes\alpha+1}=\mathrm{fin}\cdot\mathrm{fin}^{\otimes\alpha}$ and by induction hypothesis and Corollary 2.30 $\mathrm{fin}^{\otimes\alpha+1}$ is deterministic. At limit step, $\mathrm{fin}^{\otimes\alpha}=\sum_{\mathrm{fin}}\mathrm{fin}^{\otimes\alpha_{n}}$. By it is known that $\mathrm{fin}^{\alpha_{n}}$ is increasing in the Tukey order (even in the Rudin-Keisler order, see for example [3, Lemma 3.2]) and by induction hypothesis are all deterministic. Therefore, by Proposition 2.29 $\mathrm{fin}^{\otimes\alpha}$ is deterministic.
In [3] it was noticed that a generic ultrafilter for $P(X)/I$ where $I$ is a $\sigma$-ideal satisfies the $I$-p.i.p. now together with Corollary 2.27, we recover the result from [3] in our abstract settings:
Corollary 3.8.
For every $\alpha$, a generic ultrafilter $G$ for $P(\omega^{\alpha})/\mathrm{fin}^{\otimes\alpha}$ satisfies $G\cdot G\equiv_{T}G$.
Proof.
$G$ has the $\mathrm{fin}^{\otimes\alpha}$-p.i.p, and by Corollary 3.7 $\mathrm{fin}^{\otimes\alpha}$ is deterministic. Also note that $(\mathrm{fin}^{\otimes\alpha})^{\omega}\equiv_{T}\mathrm{fin}^{\otimes\alpha}$ (see [3, Thm. 3.3 & Fact 2.4]). Thus by Corollary 2.27 $G\cdot G\equiv_{T}G$. ∎
4. On ultrafilters above $I^{\omega}$
The cofinal type of $\omega^{\omega}$ came up in several early papers [28, 35] regarding the Tukey order on ultrafilters. It turns out that some general problems boil down to being Tukey above $\omega^{\omega}$. Such a problem is addressed in Theorem 0.2, which sets up the equivalence for $p$-point ultrafilter $U$, between $U\cdot U\equiv_{T}U$ and $U\geq_{T}\omega^{\omega}$. This was genralized in Theorem 0.3 which ensures that for a general ultrafilter $U$, $U\cdot U\equiv_{T}U$ is equivalent to the existence of some $I\subseteq U^{*}$ such that $U$ has the $I$-p.i.p and $U\geq_{T}I^{\omega}$. In the first part of this section, we tighten the connection between the $I$-p.i.p and being above $I^{\omega}$ for a deterministic ideal $I$. Then, in the second subsection, we shall restrict our attention to $I=\omega$.
4.1. The case of a deterministic ideal $I$
There is a slight difference between the type of equivalence in Theorem 0.2 to the equivalence $U\cdot U\equiv_{T}U$ for $p$-points and the general one in 0.3. Indeed, in the latter, the ideal $I$ can vary. Hence it is unclear in general if for a fixed $I$, the following is true: for any ultrafilter $U$ which has the $I$-p.i.p, $U\cdot U\equiv_{T}U$ iff $U\geq_{T}I^{\omega}$. Let us note first that such equivalence holds for simple ideals (and therefore also for deterministic ideals).
@@ 4.2. Ultrafilters above $\omega^{\omega}$
As observed by Dobrinen and Todorcevic, rapid ultrafilters form a subclass of those which are Tukey above $\omega^{\omega}$. By Miller [26], in the Laver model where the Borel conjecture holds181818 i.e. the model obtained by adding $\omega_{2}$-many Laver reals to a model of $CH$., there are no rapid ultrafilters. On the other hand, there are always ultrafilters which are above $\omega^{\omega}$, namely tukey-top ultrafilters. The question regarding the existence (in $ZFC$) of non-Tukey top ultrafilters is a major open problem. Hence we do not expect to have a simple $ZFC$ construction of a non-Tukey-top ultrafilter which is above $\omega^{\omega}$.
@@
Note that 00-almost-rapid is just rapid, and by $(3)$ of the previous lemma, if $\beta\leq\alpha$ then $\beta$-almost-rapid implies $\alpha$-almost-rapid. We call $U$ almost-rapid if it is $1$-almost-rapid.
@@
Let $P_{n}=\{1,...,2^{n}\}$. Let $I=\{A\subseteq\omega\mid\exists k\forall n,\ |A\cap P_{n}|\leq k\cdot n\}$. Then $I$ is a proper ideal on $\omega$. Suppose that ${\langle}P_{n}\mid n<\omega{\rangle}$ is not a counterexample for $U$ being rapid, then there is a set $X\in U$ such that $|X\cap P_{n}|\leq n$ for every $n$ and therefore $X\in I$. Hence, as long as we have $U\subseteq I^{+}$, $U$ will not be rapid. Note that
$A\in I^{+}\text{ iff for every }k,\ \text{there is }n_{k}\text{ such that }|A\cap P_{n_{k}}|>kn_{k}.$
@@
Let $V_{0}=I^{*}$. At limit steps $\delta$ we define $V_{\delta}=\bigcup_{\beta<\delta}V_{\beta}$. It is clear that $(1)-(5)$ still holds at limit steps. At successors, given $V_{\alpha}$, since we have only performed countably many steps so far, there are sets $B_{n}\in V_{\alpha}$ such that $V_{\alpha}=I^{*}[{\langle}B_{n}\mid n<\omega{\rangle}]$ where $B_{n}$ is $\supseteq$-decreasing. If either $X_{\alpha}$ or $\omega\setminus X_{\alpha}$ is already in $V_{\alpha}$, we ignore it. Otherwise, we must also have $V_{\alpha}[X_{\alpha}]\subseteq I^{+}$. If $\vec{A}_{\alpha}\not\subseteq V_{\alpha}[X_{\alpha}]$ ignore it. Otherwise, enumerate the set $\{B_{n}\cap X_{\alpha}\mid n<\omega\}\cup\{\vec{A}_{\alpha}(n)\mid n<\omega\}\subseteq V_{\alpha}[X_{\alpha}]$ by ${\langle}B^{\prime}_{n}\mid n<\omega{\rangle}$ and let $C_{n}=\cap_{m\leq n}B^{\prime}_{m}$. We apply the previous lemma to the sequence ${\langle}C_{n}\mid n<\omega{\rangle}$, and $\tau_{\alpha}$ to find $A^{*}\subseteq\omega$ such that:
@@
Since for every $n<\omega$, there is $n^{\prime}$ such that $C_{n^{\prime}}\subseteq\vec{A}_{\alpha}(n)\cap B_{n}$, $A^{*}\subseteq^{*}\vec{A}_{\alpha}(n)$, namely $A^{*}$ is a pseudo intersection of both ${\langle}B_{n}\mid n<\omega{\rangle}$ and $\vec{A}_{\alpha}$. Also, $A^{*}$ is a positive set with respect to the ideal $V_{\alpha}[X_{\alpha}]$. Otherwise, there is some $A\in I^{*}$ and $B_{n}$ such that $A^{*}\cap(A\cap B_{n}\cap X_{\alpha})=\emptyset$. But then $(A^{*}\cap B_{n}\cap X_{\alpha})\cap A=\emptyset$ which implies that $A^{*}\cap B_{n}\cap X_{\alpha}\in I$. However, $A^{*}\subseteq^{*}B_{n}\cap X_{\alpha}$, which implies that $A^{*}\in I$, contradicting property $(1)$ above in the choice of $A^{*}$. Hence we can define $V_{\alpha+1}=V_{\alpha}[X_{\alpha},A^{*}]$ and $(1)-(5)$ hold.
This concludes the recursive definition. The ultrafilter witnessing the theorem is defined by $V^{*}=\cup_{\alpha<\omega_{1}}V_{\alpha}$.
@@
The following addresses the commutativity of cofinal types between ultrafilters on different cardinals.

### v2

@@ 3. Commutativity of cofinal types
Let $\mathcal{C}$ be a commutative class, then for any two ultrafilters $U,V\in\mathcal{C}$, $Ch(U\cdot V)=Ch(V\cdot U)$.
Definition 3.7.
Let $D$ be a cofinal type, define $\mathcal{E}_{D}=\{U\mid U\cdot U\equiv_{T}U\times D\}$.
For example, $\mathcal{E}_{\{0\}}=\mathcal{E}_{\omega}=\{U\mid U\cdot U\equiv_{T}U\}$. Also $\mathcal{C}_{P}\subseteq\mathcal{E}_{\omega^{\omega}}$.
Claim 3.8.
The class $\mathcal{E}_{\{0\}}\subseteq\mathcal{E}_{\omega^{\omega}}$.
@@
Just note that $\omega^{\omega}\leq_{T}U\cdot U$ for every uniform ultrafilter $U$ and therefore if $U\cdot U\equiv_{T}U$ then $U\equiv_{T}U\times\omega^{\omega}$ and in particular $U\cdot U\equiv_{T}U\times\omega^{\omega}$. ∎
Proposition 3.9.
For every cofinal type $D$, $\mathcal{E}_{D}$ is a commutative class
Proof.
Indeed, for any $U,V\in\mathcal{E}_{D}$, by Theorem 1.7
$V\cdot U\equiv_{T}V\times U^{\omega}\equiv_{T}V\times U\cdot U\equiv_{T}V\times U\times D$
Since the formula above is symmetric for $V,U$, we conclude that $U\cdot V\equiv_{T}V\cdot U$. ∎
Proposition 3.10.
The commutative class $\mathcal{E}_{\omega^{\omega}}$ contains all Tukey-top ultrafilters, $p$-points, stable ordered union ultrafilters and generic ultrafilters for $P(\omega)/I$ where $I^{\omega}\equiv_{T}I$ is a simple $\sigma$-ideal
Proof.
Clearly Tukey top ultrafilters are in any $\mathcal{E}_{D}$, for $D$ of size at most continuum. $p$-points are in $\mathcal{E}_{\omega^{\omega}}$ as we have seen by Milovich [29], stable ordered union and generic ultrafilters for $P(\omega^{\alpha})/I$ are in fact in $\mathcal{E}_{\{0\}}$, by 2.27. ∎
@@
$\mathcal{E}_{\omega^{\omega}}$ is closed under Fubini sums.
@@
Suppose that $\{W,W_{0},W_{1},...\}\subseteq\mathcal{E}_{\omega^{\omega}}$. We need to prove that
$\sum_{W}W_{n}\cdot\sum_{W}W_{n}\equiv_{T}\sum_{W}W_{n}\times\omega^{\omega}.$
Note that $\sum_{W}W_{n}\geq_{T}\omega^{\omega}$, so we will end up getting $\sum_{W}W_{n}\cdot\sum_{W}W_{n}\equiv_{T}\sum_{W}W_{n}$. It is not hard to see that
$\sum_{W}W_{n}\cdot\sum_{W}W_{n}\geq_{T}\sum_{W}W_{n}\times\omega^{\omega}.$
For the other direction, recall that $\sum_{W}W_{n}\cdot\sum_{W}W_{n}\equiv_{T}(\sum_{W}W_{n})^{\omega}$. Let us prove that $(\sum_{W}W_{n})^{\omega}$ is a uniformly below $\mathcal{B}(W,{\langle}W_{n}\mid n<\omega{\rangle})$. Since $(\sum_{W}W_{n})^{\omega}$ is complete, Theorem 1.12 can be applied that get $\sum_{W}W_{n}\geq_{T}(\sum_{W}W_{n})^{\omega}$.
Let $f:W\times\omega^{\omega}\rightarrow W^{\omega}$ and $f_{n}:W_{n}\times\omega^{\omega}\rightarrow W_{n}^{\omega}$ be monotone and cofinal maps, which exists by the assumption that $W,W_{n}\in\mathcal{E}_{\omega^{\omega}}$. We need to define a sequence of monotone and cofinal maps ${\langle}g_{X}:W\times\prod_{n\in X}W_{n}\rightarrow(\sum_{W}W_{n})^{\omega}\mid X\in W{\rangle}$ such that $(\dagger)$ of definition 1.11 holds.
Let ${\langle}B,{\langle}A_{n}\mid n\in X{\rangle}{\rangle}\in W\times\prod_{n\in X}W_{n}$, by lemma 1.13, we may restrict ourselves to sequences satisfying that
[…]
The first step is to produce $\omega$-many functions in $\omega^{\omega}$ which are going to be the inputs of $f_{n}$’s. Fix a partition of $\omega=\biguplus_{l<\omega}Z_{l}$ such that each $Z_{l}$ is infinite. Recall that for a set $C$ on natural numbers, $C(r)$ denotes the $r^{\text{th}}$ element of $C$ in its increasing enumeration. Define $\varphi_{X,i}$ by induction on $i$. For $i=0$, $\varphi_{X,0}(k)=X(Z_{0}(k))$. Inductively, $\varphi_{X,{i+1}}(k)=\max(\varphi_{X,i}(k),X(Z_{i+1}(k)))$.
Claim 3.12.
If $Y\subseteq X$, then for every $i$, $\varphi_{X,i}\leq\varphi_{Y,i}$.
@@
Clearly, for every $m<\omega$, $X(m)\leq Y(m)$. So by definition, $\varphi_{X,0}(k)\leq\varphi_{Y,0}$. Suppose this was true for $i$, and let $k<\omega$, then by the induction hypothesis and our first observation,
$\max(\varphi_{X,i}(k),X(Z_{i+1}(k)))\geq\max(\varphi_{Y,i}(k),Y(Z_{i+1}(k)).$
The $i$’s function we will use is $h^{i}_{{\langle}A_{n}\mid n\in X{\rangle}}(k)=\min(A_{\varphi_{X,i}(k)})$. This is well defined as by definition of $\varphi_{X,i}$, $\varphi_{X,i}(k)\in X$. If $Y\subseteq X$, then by the claim $\varphi_{X,i}(k)\leq\varphi_{Y,i}(k)$, and by $(\star)$, $h^{i}_{{\langle}A_{n}\mid n\in X{\rangle}}(k)\leq h^{i}_{{\langle}A_{n}\mid n\in Y{\rangle}}(k)$. Now define $g_{X}({\langle}B,{\langle}A_{n}\mid n\in X{\rangle}{\rangle})$ by
${\langle}\pi_{X}({\langle}f(B,h^{0}_{{\langle}A_{n}\mid n\in X{\rangle}})_{m},{\langle}f_{n}(A_{n},h^{n+1}_{{\langle}A_{n}\mid n\in X{\rangle}})_{m}\mid n\in X{\rangle})\mid m<\omega{\rangle}$
The above seemingly complicated definition is nothing but the composition of the following quite natural monotone cofinal maps:
$W\times\prod_{n\in X}W_{n}\overset{(id,{\langle}h^{i}_{*}\mid i<\omega{\rangle})}{\longrightarrow}W\times\prod_{n\in X}W_{n}\times(\omega^{\omega})^{\omega}\rightarrow(W\times\omega^{\omega})\times\prod_{n\in X}(W_{n}\times\omega^{\omega})\overset{(f,{\langle}f_{n}\mid n\in X{\rangle})}{\longrightarrow}$
$\overset{(f,{\langle}f_{n}\mid n\in X{\rangle})}{\longrightarrow}W^{\omega}\times\prod_{n\in X}W_{n}^{\omega}\longrightarrow(W\times\prod_{n\in X}W_{n})^{\omega}\overset{\pi_{X}^{\omega}}{\longrightarrow}(\sum_{W}W_{n})^{\omega}$
So $g_{X}$ is clearly monotone cofinal as the composition of such functions. To see $(\dagger)$, let $Y\subseteq X$, we ensured that $h^{i}_{{\langle}A_{n}\mid n\in X{\rangle}}\leq h^{i}_{{\langle}A_{n}\mid n\in Y{\rangle}}$. Since $f,f_{n}$ are monotone functions, for any $m<\omega$, and any $n\in Y$,
$f(B,h^{0}_{{\langle}A_{n}\mid n\in Y{\rangle}})_{m}\subseteq f(B,h^{0}_{{\langle}A_{n}\mid n\in X{\rangle}})_{m}$
and
$f_{n}(A_{n},h^{n+1}_{{\langle}A_{n}\mid n\in Y{\rangle}})_{m}\subseteq f_{n}(A_{n},h^{n+1}_{{\langle}A_{n}\mid n\in X{\rangle}})_{m}.$
By definition of $\pi_{X}$ and $\pi_{Y}$, as unions of the relevant sets (see p. 9) we conclude that $g_{Y}(\pi_{X,Y}({\langle}B,{\langle}A_{n}\mid n\in X{\rangle}{\rangle}))\geq g_{X}({\langle}B,{\langle}A_{n}\mid n\in X{\rangle}{\rangle})$, as wanted. ∎
Given a cofinal type $D$, by a (yet) finer handling of coherent sequences of functions, we can define an analog of condition $(\dagger)$, that ensure we can amalgamate monotone cofinal maps $g_{X}:U\times\prod_{n\in X}V_{n}\times D\rightarrow\mathbb{P}$, to a monotone cofinal map $g:\sum_{U}V_{n}\times D\rightarrow\mathbb{P}$. Thus proving an analog of Theorem 1.12 that $\sum_{U}V_{n}\times D$ is the greatest lower bound of $\{U\times\prod_{n\in X}V_{n}\times D\mid X\in U\}$ among uniform lower bounds. Then the proof of the previous Theorem generalizes to give the next theorem. Since the proofs are essentially the same, we do not include them.
Theorem 3.13.
Suppose that $D^{\omega}\equiv_{T}D$ then $\mathcal{E}_{D}$ is closed under Fubini sums.
We do not know whether the class $\mathcal{E}_{\omega^{\omega}}$ is provably all uniform ultrafilters on $\omega$. We conjecture that consistently it is not but this will require to produce new examples of non-Tukey top ultrafilters, as it is unknown whether there is a basically generated ultrafilter which is not in the (transfinite) closure of the $p$-points under Fibini sums, and the only examples for non-Tukey-top non-basically generated ultrafilters are generic ultrafilters for $P(\omega^{\otimes\alpha})/\mathrm{fin}^{\otimes\alpha}$.
4. On ultrafilters above $I^{\omega}$
The cofinal type of $\omega^{\omega}$ came up in several papers [28, 36] regarding the Tukey order on general ultrafilters. Milovich asked whether there is an ultrafilter $U$ such that $(U,\supseteq)$ is Tukey equivalent to $\omega^{\omega}$. Let us point out that a negative answer is a straightforward corollary [^20] Milovich’s question appeared only 4 years after Solecki and Todorcevic’s result. of Theorem 0.3:
Proposition 4.1.
There is no non-principal ultrafilter $U$ over $\omega$ such that $(U,\supseteq)\equiv_{T}\omega^{\omega}$.
@@ 4. On ultrafilters above $I^{\omega}$
By Sierpinski [35], a non-principal ultrafilter over $\omega$ is a non-measurable set as a subset of $2^{\omega}$ and in particular non-analytic. An ultrafilter $U$ with the topology inherited from $2^{\omega}$ is a separable metric space and the set of predecessors is compact and $\omega^{\omega}$ is a basic analytic order, hence by Theorem 0.3, $(U,\supseteq)\not\leq_{T}\omega^{\omega}$. ∎
It turns out that some general problems boil down to being Tukey above $\omega^{\omega}$. Such a problem is addressed in Theorem 0.4, which sets up the equivalence for $p$-point ultrafilter $U$, between $U\cdot U\equiv_{T}U$ and $U\geq_{T}\omega^{\omega}$. This was generalized in Theorem 0.5 which ensures that for a general ultrafilter $U$, $U\cdot U\equiv_{T}U$ is equivalent to the existence of some $I\subseteq U^{*}$ such that $U$ has the $I$-p.i.p and $U\geq_{T}I^{\omega}$. In the first part of this section, we tighten the connection between the $I$-p.i.p and being above $I^{\omega}$ for a deterministic ideal $I$. Then, in the second subsection, we shall restrict our attention to $I=\omega$.
4.1. The case of a deterministic ideal II
There is a slight difference between the type of equivalence in Theorem 0.4 to the equivalence $U\cdot U\equiv_{T}U$ for $p$-points and the general one in 0.5. Indeed, in the latter, the ideal $I$ can vary. Hence it is unclear in general if for a fixed $I$, the following is true: for any ultrafilter $U$ which has the $I$-p.i.p, $U\cdot U\equiv_{T}U$ iff $U\geq_{T}I^{\omega}$. Let us note first that such equivalence holds for simple ideals (and therefore also for deterministic ideals).
@@ 4.1. The case of a deterministic ideal II
The previous proposition is somewhat different from [3, Thm. 1.18], as here $I$ is a fix simple ideal.
@@ 4.2. Ultrafilters above $\omega^{\omega}$
As observed by Dobrinen and Todorcevic, rapid ultrafilters form a subclass of those which are Tukey above $\omega^{\omega}$. By Miller [26], in the Laver model where the Borel conjecture holds212121i.e. the model obtained by adding $\omega_{2}$-many Laver reals to a model of $CH$., there are no rapid ultrafilters. On the other hand, there are always ultrafilters which are above $\omega^{\omega}$, namely tukey-top ultrafilters. The question regarding the existence (in $ZFC$) of non-Tukey top ultrafilters is a major open problem. Hence we do not expect to have a simple $ZFC$ construction of a non-Tukey-top ultrafilter which is above $\omega^{\omega}$.
@@
Note that 0-almost-rapid is just rapid, and by $(3)$ of the previous lemma, if $\beta\leq\alpha$ then $\beta$-almost-rapid implies $\alpha$-almost-rapid. We call $U$ almost-rapid if it is $1$-almost-rapid.
@@
Let $P_{n}=\{1,...,2^{n}\}$. Let $I=\{A\subseteq\omega\mid\exists k\forall n,\ |A\cap P_{n}|\leq k\cdot n\}$. Then $I$ is a proper ideal on $\omega$. Suppose that ${\langle}P_{n}\mid n<\omega{\rangle}$ is not a counterexample for $U$ being rapid, then there is a set $X\in U$ such that $|X\cap P_{n}|\leq n$ for every $n$ and therefore $X\in I$. Hence, as long as we have $\subseteq I^{+}, UU$ will not be rapid. Note that
[…]
@@
Let $V_{0}=I^{*}$. At limit steps $\delta$ we define $V_{\delta}=\bigcup_{\beta<\delta}V_{\beta}$. It is clear that (1)(5)(1)-(5) still holds at limit steps. At successors, given $V_{\alpha}$, since we have only performed countably many steps so far, there are sets $B_{n}\in V_{\alpha}$ such that $V_{\alpha}=I^{*}[{\langle}B_{n}\mid n<\omega{\rangle}]$ where $B_{n}$ is $\supseteq$-decreasing. If either $X_{\alpha}$ or $\omega\setminus X_{\alpha}$ is already in $V_{\alpha}$, we ignore it. Otherwise, we must also have $V_{\alpha}[X_{\alpha}]\subseteq I^{+}$. If $\vec{A}_{\alpha}\not\subseteq V_{\alpha}[X_{\alpha}]$ ignore it. Otherwise, enumerate the set
$\{B_{n}\cap X_{\alpha}\mid n<\omega\}\cup\{\vec{A}_{\alpha}(n)\mid n<\omega\}\subseteq V_{\alpha}[X_{\alpha}]$
by ${\langle}B^{\prime}_{n}\mid n<\omega{\rangle}$ and let $C_{n}=\cap_{m\leq n}B^{\prime}_{m}$. We apply the previous lemma to the sequence ${\langle}C_{n}\mid n<\omega{\rangle}$, and $\tau_{\alpha}$ to find $A^{*}\subseteq\omega$ such that:
@@
Since for every $n<\omega$, there is $n^{\prime}$ such that $C_{n^{\prime}}\subseteq\vec{A}_{\alpha}(n)\cap B_{n}$, $A^{*}\subseteq^{*}\vec{A}_{\alpha}(n)$, namely $A^{*}$ is a pseudo intersection of both ${\langle}B_{n}\mid n<\omega{\rangle}$ and $\vec{A}_{\alpha}$. Also, $A^{*}$ is a positive set with respect to the ideal $V_{\alpha}[X_{\alpha}]$. Otherwise, there is some $A\in I^{*}$ and $B_{n}$ such that $A^{*}\cap(A\cap B_{n}\cap X_{\alpha})=\emptyset$. But then $(A^{*}\cap B_{n}\cap X_{\alpha})\cap A=\emptyset$ which implies that $A^{*}\cap B_{n}\cap X_{\alpha}\in I$. However, $A^{*}\subseteq^{*}B_{n}\cap X_{\alpha}$, which implies that $A^{*}\in I$, contradicting property $(1)$ above in the choice of $A^{*}$. Hence we can define $V_{\alpha+1}=V_{\alpha}[X_{\alpha},A^{*}]$ and (1)(5)(1)-(5) hold.
This concludes the recursive definition. The ultrafilter witnessing the theorem is defined by $V^{*}=\bigcup_{\alpha<\omega_{1}}V_{\alpha}$.
@@
The following addresses the possibility of full commutativity of cofinal types between ultrafilters on $\omega$ and on different cardinals.
@@
Is it provable that for any two ultrafilters $U,V$, $U\cdot V\equiv_{T}V\cdot U$?
Question 5.9.
@@
the results in § 4 suggest that the Tukey-type of $I^{\omega}$ as an important role in the analysis of non Tukey-top ultrafilters. Note that the same argument which worked for $\omega^{\omega}$ in Proposition 4.1 works for $I^{\omega}$, hence leading to the conclusion that if $I$ is an analytic ideal the no non-principal ultrafilter can be Tukey equivalent to $I^{\omega}$. The following question is more open-ended:
Question 5.10.
How can we force different values of the Tukey-type of $I^{\omega}$, when $I\not\equiv_{T}\omega^{\omega}$?